OpenStudy (anonymous):

Prove that the relation is an equivalence relation. The relation T on RxR given by (x,y)T(a,b) iff x^2+y^2=a^2+b^2. Sketch the equivalence class of (1,2)

OpenStudy (akashdeepdeb):

Can you prove Reflexivity?

OpenStudy (anonymous):

I'm not sure how to even begin this.

OpenStudy (ikram002p):

i think u have to show that T is equivalence relation mm 1- the reflexive law holds . 2- the symmetric law holds . 3- the transitive law holds .

OpenStudy (ikram002p):

and as i see for this (x,y)T(a,b) iff x^2+y^2=a^2+b^2. its a circl x^2+y^2=r^2=a^2+b^2. so for (1,2) its a circle with radues \(\sqrt 5\) center 0,0